Suppose we have a metal with a work function of 4.46 eV,

What is the maximum speed, in meters per second, of electrons ejected from this metal by photons of light with wavelength 85mn?

Vmax=?
Answer #1

Energy of photon = hc / lambda

E = (1240 eV nm ) / 85 nm = 14.59 eV

using Energy of photon = Work function + energy of released electron

14.59 = 4.46 + Ee

Ee = 10.13 eV


and energy = m v^2 /2

(9.109 x 10^-31 kg ) v^2 /2   = ( 10.13 x 1.6 x 10^-19 J )

v = 1.89 x 10^6 m/s